OpenStudy (anonymous):

lim as x goes to 0: (tan^2(4x))/((2x)^2)

OpenStudy (anonymous):

lol, I hope you found out how to do these without L'Hopital's rule Again, I would advise using it, it's super effective I'll solve it if you need it

OpenStudy (anonymous):

haha, i am substituting. Goal later on is to learn hopital!

OpenStudy (anonymous):

lol good :) It shouldn't be too hard, you'll do fine glgl

OpenStudy (anonymous):

If you don't know l'hopital you can still do this one with the rule about sinx/x.

OpenStudy (anonymous):

got up to 4/cos^2(4x)

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

Where'd your sins go?

OpenStudy (anonymous):

they became 1's

OpenStudy (anonymous):

figured it out, plug in the 0 cause cos of 0 =1!

OpenStudy (anonymous):

Fair enough. Then I think the only problem is that the 4 should be on bottom.

OpenStudy (anonymous):

answer is 4

OpenStudy (anonymous):

But yeah, plug in for the cos

OpenStudy (anonymous):

thanks everyone!

OpenStudy (anonymous):

No, it's 1/4.

OpenStudy (anonymous):

\[{tan^2(4x)\over(2x)^2} = ({sin(4x)\over x})^2({1 \over 4 cos^2(4x)})\] \[\implies \lim_{x\rightarrow 0} {tan^2(4x)\over(2x)^2} = \lim_{x\rightarrow 0} ({sin(4x)\over x})^2({1 \over 4 cos^2(4x)}) = {1\over 4}\]

OpenStudy (anonymous):

no it's 4

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

Then you typed something wrong in the original equation.

OpenStudy (anonymous):

teacher said so

OpenStudy (anonymous):

no, you got it written correctly

OpenStudy (anonymous):

You have a 4 in the denominator. Where did it go?

OpenStudy (anonymous):

one sec polpak; I am asking the teacher. I think you might be right>

OpenStudy (anonymous):

Yes I am.

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

going to have to wait until the lecture is done _ can I send it to you _ i am really interested to see the result

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thanks a ton for your help!